OpenStudy (anonymous):

All of the following are equal EXCEPT _____. $360.00 15% off $420.00 10% off $400.00 25% off $480.00 giving a metal

OpenStudy (mathmate):

hint: the calculator is your best friend!

OpenStudy (anonymous):

i dont use calculators but i guess so @mathmate

OpenStudy (mathmate):

You're quite right. The above sums can be all handled mentally, although not expected of most students.

OpenStudy (anonymous):

i guess

OpenStudy (anonymous):

1) 10% of 400 (or 0.1 x 400) = 40. 400 - 40 = 360. It's not this one. 2) 25% of 480 (or o.25 x 480) = 120. 480 - 120 = 306. It's not this one. 3) 15% of 420 = 63. 420 - 63 = 357. THEREFORE - the odd one out is 15% off of 420.

OpenStudy (anonymous):

thank you @jakelipman